solve each system using the substitutions method.
x=y-5
2x+5y=4

Answer by mananth(16946)

x=y-5
2x+5y=4
2(y-5) +5y=4
2y-10+5y=4
7y=14
y=2
x=y-5
x=2-5
x=-3
Ananth